OpenStudy (anonymous):

Write the normal form of the equation: 9x-2y+4=0

OpenStudy (goldphenoix):

if you mean set the equation so it equals x it is \[x = \frac{ 2y-4 }{ 9 }\]

OpenStudy (anonymous):

If you are finding y=..then 9x-2y+4=0 +2y = 2y (9x+4=2y)/2 4.5x+2=y y=4.5x+2 i think thats correct